# The scalar-torsion gravity corrections in the first-order inflationary models

### v1

@@ IV State parameter at the end of the inflationary stage
where $\tilde{p}$ and $\tilde{\rho}$ are the pressure and energy density of a scalar field correspondingly.
At the inflationary stage, due to condition $\tilde{\epsilon}\ll 1$ from expressions (36) one has $\tilde{w}\simeq-1$. The end of the accelerated expansion stage of the early universe is defined by condition $\tilde{\epsilon}=1$ with corresponding state parameter $\tilde{w}_{E}=-1/3$ which follows from (36) for the case of the minimal coupling [1, 2].
For the case of non-minimal coupling $n\neq 0$, the state parameter is
@@
Taking to account expression (32), from expression (37) we obtain
@@
At the inflationary stage, due to conditions $\tilde{\epsilon}\ll 1$ and $\epsilon\ll 1$ from expressions (36) and (38) one has $\tilde{w}\simeq-1$ and $w\simeq-1$.
For the special case $n=2$, due to expression (38), at the end of inflation $\epsilon=1$ the state parameter is $w_{E}=1$ for any value of the constant parameter $\lambda$. This case corresponds to the post-inflationary stage of kinetic energy domination (kination), which precedes the further stage of radiation domination [37, 38, 39].
For the case $n\neq 2$, at the end of inflation $\epsilon=1$ from condition $w_{E}=-1/3$ and expression (38), we obtain following relation
@@
Also, from condition $\lambda^{n}>0$, we obtain following restrictions on the parametrization constant $-1<n<2$ for this case.
Thus, from equations (30)–(31) and (39), we obtain following relations
@@
To determine what type of completion of the inflationary stage is implemented for the models under consideration, it is necessary to determine the upper limit on the parametrization constant $n$ in addition to the lower limit (26). This estimate can be obtained on the base of the analysis of the influence of non-minimal coupling between a scalar field and torsion on the parameters of the cosmological perturbations in the inflationary models under consideration.
V Adjusted parameters of cosmological perturbations
From the standpoint of the inflationary paradigm, the origin of the large-scale structure of the universe is due to quantum fluctuations of the scalar field and corresponding scalar perturbations of the space-time metric at the inflationary stage of the evolution of the universe. Also, the theory of cosmological perturbations predicts the presence of relict gravitational waves or tensor perturbations in the present era which were also induced at the inflationary stage [1, 2].
The influence of cosmological perturbations on the anisotropy and polarization of CMB makes it possible to determine the correctness of inflationary models based on observational constraints on the values of the parameters of cosmological perturbations [1, 2]. According to modern observational data [40, 41], constraints on the parameters of cosmological perturbations are
@@ V Adjusted parameters of cosmological perturbations
where $P_{S}$ is the power spectrum of scalar perturbations, $n_{S}$ is the spectral index of scalar perturbations and $r$ is the tensor-to-scalar ratio.
It is obvious that the non-minimal coupling of the scalar field and torsion affects the evolution of cosmological perturbations and induces some corrections to the values of the parameters of cosmological perturbations relative to the case of minimal coupling. The analysis of the evolution of cosmological disturbances in inflationary models based on the TEGR and its modifications was considered earlier in Refs. [15, 16, 17, 18, 19].
In Ref. [18] cosmological perturbations for the inflationary models based on the generalized scalar-torsion gravity were considered based on the Arnowitt-Deser-Misner (ADM) decomposition of the tetrad field to define the perturbed metrics. By this approach, the parameters of cosmological perturbations at the crossing of the Hubble radius $k=aH$ can be written as follows [18]
@@
${\mathcal{P}}_{T}=\frac{H^{2}}{2\pi^{2}Q_{T}}=\frac{2H^{2}}{\pi^{2}F},$
@@
$n_{T}=-2\epsilon-\delta_{F},$
$r=\frac{{\mathcal{P}}_{T}}{{\mathcal{P}}_{S}}=16c_{s}\epsilon_{s},$
@@
the slow-roll parameters are
@@
and the velocity of the scalar perturbations is
@@
Also, we note, that all slow-roll parameters satisfy the condition $\epsilon,\delta,\delta_{F},\epsilon_{s},\eta_{s},s\ll 1$ during inflation, and condition $\dot{H}<0$ implies $\epsilon>0$ as well.
For the inflationary models based on the action (1) functions $w_{1}$, $w_{2}$, $w_{3}$ can be defined as [18]
@@
As one can see, in the general case, the velocity of propagation of scalar perturbations is a function of cosmic time $c_{S}=c_{S}(t)$. Thus, in the general case, corrections induced by the non-minimal coupling of the scalar field and torsion must be considered for each specific model separately. Nevertheless, for the power-law connection between the non-minimal coupling function and the Hubble parameter $F=(H/\lambda)^{n}$ from (53)–(V), we obtain the following expression for the velocity of propagation of scalar perturbations
$c_{S}^{2}=\frac{1-n}{1+n}=const,$ (56)
for an arbitrary model’s parameters.
The condition $c_{S}^{2}>0$ for expression (56) leads to the following conditions on the parametrization constant
@@
A super-luminal propagation cosmological perturbations is possible in the presence of the time-dependent homogeneous scalar field [42, 43, 44] and for different modified gravity theories [45, 46, 47] as well. Also, we note, that due to constraint (57), the possible post-inflationary kination stage (for $n=2$) can not be realized in these cosmological models.
@@
The difference in the propagation velocities of scalar and tensor perturbations corresponds to the fact that they cross the Hubble radius with different wavelengths, namely
@@
unlike the case of the minimal coupling $n=0$ for which $c_{S}=c_{T}=1$ and the wavelengths of scalar and tensor perturbations are equal $\lambda_{S}=\lambda_{T}$.
Also, for the case $F=(H/\lambda)^{n}$, from (50) we obtain
@@
The conditions $Q_{S}>0$ and $Q_{T}>0$ mean that there are no Laplacian instabilities and ghosts in these models [18].
The slow-roll parameters (51)–(52) for the case $F=(H/\lambda)^{n}$ are related as follows
@@
Thus, parameters of cosmological perturbations (46)–(49) can be written as
@@
${\mathcal{P}}_{T}=\frac{2\lambda^{n}}{\pi^{2}}H^{2-n}_{\ast},$ (65)
$n_{S}-1=(n-4)\epsilon_{\ast}+2\delta_{\ast},$ (66)
$r=16(1-n)\sqrt{\frac{1-n}{1+n}}\,\epsilon_{\ast},$ (67)
$n_{T}=-(2-n)\epsilon_{\ast}.$ (68)
For the case of the minimal coupling $n=0$ we obtain
@@
We note, that all parameters of cosmological perturbations are considered at the crossing of the Hubble radius ($k=aH$ and $\tilde{k}=\tilde{a}h$ for some cosmic times $\tilde{t}=\tilde{t}_{\ast}$ and $t=t_{\ast}$, where $t_{\ast}\neq\tilde{t}_{\ast}$ and, correspondingly, $H^{n+1}_{\ast}\neq h_{\ast}$).
Since in both cases (non-minimal and minimal coupling) cosmological models can satisfy condition (42), the relation between the cosmic times of the crossing of the Hubble radius $\xi=\left(t_{\ast}/\tilde{t}_{\ast}\right)$ can be obtained from the equation for certain inflationary models
@@
Since condition (42) can always be satisfied by choosing the time of crossing the Hubble radius or model’s parameters, we consider the spectral index of scalar perturbations and the tensor-scalar ratio as the main parameters necessary for verifying inflationary models by observational constraints (43)–(44). Also, in accordance with restrictions (57), condition $\epsilon_{\ast}>0$ and expression (68), on has red tilt tensor spectrum $n_{T}<0$ for arbitrary cosmological models under consideration.
Also, we note, that expressions for the slow-roll parameters (32)–(35) leads to the following relation
@@
Since relation (75) must be satisfied for any cosmic times, we get following conditions
$\tilde{\delta}/\tilde{\epsilon}=const,$ (76)
$\delta/\epsilon=const,$ (77)
which restrict the possible types of inflationary models.
VI Inflationary models with linear dependence $r\sim(1-n_{S})$
Now we will consider the dependence of the tensor-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$. Since the value of the spectral index of scalar perturbations is $n_{S}\simeq 0.97$ and $1-n_{S}\simeq 0.03\ll 1$, we can write the dependence $r=r(1-n_{S})$ as follows

### v3

@@ IV Equation of state (EoS) parameter at the end of the inflationary stage
where $\tilde{p}$ and $\tilde{\rho}$ are respectively the pressure and energy density of a scalar field.
At the inflationary stage, since $\tilde{\epsilon}\ll 1$ one can interpret from Eq. (43), $\tilde{w}\simeq-1$. The end of the accelerated expansion stage of the early Universe is defined by condition $\tilde{\epsilon}=1$ with corresponding EoS parameter $\tilde{w}_{E}=-1/3$ which follows from (43) for the case of the minimal coupling [1, 2].
For the case of non-minimal coupling ($n\neq 0$), the EoS parameter is
@@
Using Eq. (39) in Eq. (44), one can get
@@
At the inflationary stage, since $\tilde{\epsilon}\ll 1$ and $\epsilon\ll 1$ from expressions (43) and (45), one can find $\tilde{w}\simeq-1$ and $w\simeq-1$.
For the special case $n=2$, due to Eq. (45), at the end of inflation $\epsilon=1$ the EoS parameter is $w_{E}=1$ for any value of the constant parameter $\lambda$. This case corresponds to the post-inflationary stage of kinetic energy domination (kination), which precedes the further stage of radiation domination [64, 65, 66].
For the case $n\neq 2$, at the end of inflation $\epsilon=1$ from condition $w_{E}=-1/3$ and Eq. (45), we obtain following relation
@@
Again, from condition $\lambda^{n}>0$, we have the restrictions on the parametrization constant $-1<n<2$. Now, from Eqs. (37)–(38) and Eq. (46), we obtain the following relations
@@
To determine what type of completion of the inflationary stage is implemented for the models under consideration, it is necessary to determine the upper limit on the parametrization constant $n$ in addition to the lower limit (33). This estimate can be obtained based on the analysis of the influence of non-minimal coupling between a scalar field and torsion on the parameters of the cosmological perturbations in the inflationary models.
V Cosmological perturbations
From the standpoint of inflationary paradigm, the origin of the large-scale structure of the Universe is due to quantum fluctuations of the scalar field and corresponding scalar perturbations of the space-time metric at the inflationary stage. Also, the theory of cosmological perturbations predicts the presence of relict gravitational waves or tensor perturbations in the present era which were also induced at the inflationary stage [1, 2].
The influence of cosmological perturbations on anisotropy and polarization of CMB makes it possible to determine the correctness of inflationary models based on observational constraints on the values of the parameters of cosmological perturbations [1, 2]. Current observational data [67, 68] suggest the constraints on the cosmological perturbation parameters as,
@@ V Cosmological perturbations
where $P_{S}$ represents the power spectrum of scalar perturbations, $n_{S}$ and $r$ respectively be the spectral index of scalar perturbations and tensor-to-scalar ratio.
It is obvious that the non-minimal coupling of the scalar field and torsion affects the evolution of cosmological perturbations and induces some corrections to the values of the parameters of cosmological perturbations relative to the case of minimal coupling. The analysis of the evolution of cosmological perturbations in inflationary models based on TEGR and its modifications was studied earlier in Refs. [30, 31, 32, 33, 34]. In generalized scalar-torsion gravity, to define the perturbed metric, the cosmological perturbations for the inflationary models were studied based on Arnowitt-Deser-Misner (ADM) decomposition of the tetrad field [33]. The parameters of cosmological perturbations at the crossing of the Hubble radius $k=aH$ can be written as,
@@
${\mathcal{P}}_{T}=\frac{H^{2}}{2\pi^{2}Q_{T}}=\frac{2H^{2}}{\pi^{2}F},$ (53)
@@
$n_{T}=-2\epsilon-\delta_{F},$ (55)
$r=\frac{{\mathcal{P}}_{T}}{{\mathcal{P}}_{S}}=16c_{s}\epsilon_{s},$ (56)
@@
and the slow-roll parameters are
@@
The velocity of the scalar perturbations becomes,
@@
Here the slow-roll parameters satisfy $\epsilon\ll 1,\delta\ll 1,\delta_{F}\ll 1,\epsilon_{s}\ll 1,\eta_{s}\ll 1,s\ll 1$ during inflation. Also for $\dot{H}<0$, we can get $\epsilon>0$.
In the inflationary models for action (1), the functions $w_{1}$, $w_{2}$, $w_{3}$ can be defined as [33],
@@
We also note that in the case of scalar-torsion gravity violation of local Lorentz symmetry induces additional gauge degrees of freedom, which arises as a result of the different choice of the particular perturbed tetrad frame. These gauge degrees of freedom correspond to the additional scalar mode, the transverse vector mode, and the (pseudo) vector mode [33, 34]. This property of cosmological inflation models based on scalar-torsion gravity was discussed earlier, for example, in [30, 31, 32, 33, 34].
However, in [33] it was shown that for the inflationary models based on the action (7) Lorentz violating stage with essential contribution to inflation occurs deep inside the horizon, and these additional modes quickly decay with expansion of the early Universe. Also, in this case, the power spectrum of the scalar perturbations is calculated as the two-point correlation function based on the standard second-order action implying negligible non-Gaussianity [33]. Thus, in accordance with [33] we will consider standard scalar and tensor modes in the linear order of the cosmological perturbations theory at the crossing of the Hubble radius $(k=aH)$ for inflationary models under consideration.
V.1 Velocities of scalar and tensor perturbations
In general case, the velocity of propagation of scalar perturbations is a function of cosmic time [$c_{S}=c_{S}(t)$]. So, the corrections induced by the non-minimal coupling of scalar field and torsion must be considered for each specific model separately. Nevertheless, for the power-law connection between the non-minimal coupling function and the Hubble parameter $F=(H/\lambda)^{n}$ from Eqs. (60)–(V), we obtain the following expression for the velocity of propagation of scalar perturbations
$c_{S}^{2}=\frac{1-n}{1+n}=constant,$ (63)
for an arbitrary parameter of the model.
For $c_{S}^{2}>0$, Eq. (63) leads to the following constraint on the constant,
@@ V.1 Velocities of scalar and tensor perturbations
Fig.1 shows the dependence of the propagation speed of scalar perturbations on the parametrization constant $c_{S}=c_{S}(n)$. This dependence demonstrates that, in the general case, the models of cosmological inflation under consideration imply three different regimes of propagation of scalar perturbations: sub-luminal regime, luminal regime and super-luminal regime.
A super-luminal propagation cosmological perturbations is possible in the presence of the time-dependent homogeneous scalar field [69, 70, 71] and for different modified gravity theories [72, 73, 74] as well. Also, we note that due to constraint (64), the possible post-inflationary kination stage (for $n=2$) can not be realized in these cosmological models.
@@
The difference in propagation velocities of scalar and tensor perturbations leads to the fact that the Hubble radius was crossed by them in different wavelengths, namely
@@
Whereas, in case of the minimal coupling for which $c_{S}=c_{T}=1$, the wavelengths of scalar and tensor perturbations are equal ($\lambda_{S}=\lambda_{T}$).
On the basis of expressions (46) and (63) we can define constant model’s parameters as follows
$n=\frac{1-c_{S}^{2}}{1+c_{S}^{2}},$ (68)
$\lambda^{n}=\frac{1}{4}\left(1+3c_{S}^{2}\right).$ (69)
For the case of the minimal coupling with velocity of the scalar perturbations $c_{S}^{2}=1$ one has $n=0$ and $\lambda=1$.
Therefore, initial power-law connection (18) between coupling function and the Hubble parameter can be redefined in terms of the velocity of propagation of scalar perturbations as
$F(\phi(t))=\left(\frac{4}{1+3c_{S}^{2}}\right)\left(H(t)\right)^{\frac{1-c_{S}^{2}}{1+c_{S}^{2}}}.$ (70)
Thus, the influence of non-minimal coupling between scalar field and torsion on the cosmological parameters can be parameterized by the speed of the scalar perturbations. However, for convenience of presenting the results we will use parameters $n$ and $\lambda$ taking into account relations (68)–(69).
V.2 Parameters of cosmological perturbations
For the case $F=(H/\lambda)^{n}$, from Eq. (57), we obtain
@@ V.2 Parameters of cosmological perturbations
The conditions $Q_{S}>0$ and $Q_{T}>0$ indicate that there are no Laplacian instabilities and ghosts in these models [33]. Also, the stability of scalar perturbations is defined by condition $c^{2}_{S}>0$ corresponding to constraint (64).
From (63) and (V.2) we get
$Q_{S}=\left(\frac{2}{1+c^{2}_{S}}\right)\left(\frac{H}{\lambda}\right)^{n}\epsilon>0.$ (73)
Thus, in this case, scalar perturbations are stable in the super-luminal regime as well. However, it will be shown below that the super-luminal regime of propagation of scalar perturbations cannot be realized in these models.
For $F=(H/\lambda)^{n}$, the slow-roll parameters Eqs. (58)–(59) can be related as,
@@
Thus, the cosmological perturbations (52)–(56) parameters can be written as
@@
${\mathcal{P}}_{T}=\frac{2\lambda^{n}}{\pi^{2}}H^{2-n}_{\ast},$
$n_{S}-1=(n-4)\epsilon_{\ast}+2\delta_{\ast},$
$r=16(1-n)\sqrt{\frac{1-n}{1+n}}\,\epsilon_{\ast},$
$n_{T}=-(2-n)\epsilon_{\ast}.$
For the case of the minimal coupling, we obtain
@@
All parameters of cosmological perturbations are considered at the crossing of the Hubble radius ($k=aH$ and $\tilde{k}=\tilde{a}h$ for some cosmic times $\tilde{t}=\tilde{t}_{\ast}$ and $t=t_{\ast}$, where $t_{\ast}\neq\tilde{t}_{\ast}$ and, correspondingly, $H^{n+1}_{\ast}\neq h_{\ast}$).
Since both minimal and non-minimal coupling cosmological models satisfy condition (49), the relation between the cosmic times of the crossing of the Hubble radius $\xi=\left(t_{\ast}/\tilde{t}_{\ast}\right)$ can be obtained from the equation for certain inflationary models
@@
Since condition (49) can always be satisfied by choosing the time of crossing the Hubble radius or model’s parameters, the spectral index of scalar perturbations and the tensor-scalar ratio as the main parameters necessary for verifying inflationary models by observational constraints (50)–(51). Also, in accordance with restrictions (64), condition $\epsilon_{\ast}>0$ and expression (79), on has red tilt tensor spectrum $n_{T}<0$ for arbitrary cosmological models under consideration.
V.3 Transition to inflation based on $f(T)$-gravity
Now let us consider the possibility of reducing the model based on action (1) to the case of $f(T)$-gravity. This consideration is motivated by the fact that $f(T)$-gravity model includes the super-luminal mode that could arise from one of the extra degrees of freedom as shown in [75, 76]. Thus, we will testify the possibility to obtain the additional gauge scalar modes by transition to $f(T)$-gravity in the inflationary models under consideration.
Under the condition of zero kinetic energy of a scalar field $\omega=0$ and for relations
$F(\phi)=\phi=\frac{df(T)}{dT},$ (86)
$V(\phi)=f(T(\phi))-\frac{\phi}{2}T(\phi),$ (87)
action (7) is reduced to the following form
$S=\int d^{4}xef(T),$ (88)
corresponding to effective $f(T)$-gravity model.
Also, using expressions (7), (10) and (18) under conditions $\omega=0$ and $V=0$ we can obtain partial case of $f(T)$-gravity model, namely
$S\sim\int d^{4}xeT^{n/2}.$ (89)
However, from expressions (57) and (61)–(V) for $\omega=0$ one has $Q_{S}=0$. This result corresponds to the absence of scalar perturbations. Also, from (V.2) and $Q_{S}=0$ we can conclude that condition (18) implies that cosmological dynamics for $f(T)$-gravity corresponds to the pure de Sitter stage with $H=constant$ in the models under consideration. As one can see from relation (10), for constant Hubble parameter $f(T)$-gravity degenerates to the case $f(T)=constant$ in FLRW space-time without scalar field.
Thus, to construct correct inflationary models with quasi de Sitter dynamics and non-zero scalar perturbations we will consider condition $\omega\neq 0$.
VI Condition of the exit from inflation
Expressions for the slow-roll parameters (39)–(42) leads to the following relation
@@ VI Condition of the exit from inflation
Since relation (90) must be satisfied for any cosmic time, we get the following conditions
$\tilde{\delta}/\tilde{\epsilon}=constant,$ (91)
$\delta/\epsilon=constant,$ (92)
which restricts the possible types of inflationary models.
Based on the definitions of the slow-roll parameters (39)–(40) and the relation (92), we get following types of inflationary dynamics:
1. Intermediate inflation with following Hubble parameter and scale factor
$H(t)=A\gamma t^{\gamma-1},$ (93)
$a(t)=a_{s}\exp(At^{\gamma}),$ (94)
where $a_{s}$ is the scale factor at the beginning of inflation, $A>0$ and $10<\gamma<$ 1.
Intermediate inflation was considered earlier in [77, 78, 79, 80, 81, 82]. Intermediate inflation based on $f(T)$-gravity was also considered in [81, 82, 80].
Slow-roll parameters (39)–(40) corresponding to Hubble parameter (93) are
$\epsilon=\frac{(1-\gamma)}{A\gamma t^{\gamma}},$ (95)
$\delta=\frac{(2-\gamma)}{2A\gamma t^{\gamma}}.$ (96)
Since slow-roll parameter $\epsilon$ is a decreasing function, it is impossible to simultaneously fulfill the slow-roll condition $\epsilon\ll 1$ during inflation and exit from inflation $\epsilon_{e}=\epsilon(t=t_{e})=1$. In this case, $t_{e}$ is the time of the end of inflation, which is defined from condition $\ddot{a}=0$ or $\epsilon=1$, since
$\ddot{a}=aH^{2}(1-\epsilon).$ (97)
Thus, there is no exit from inflation by violation of the slow-roll conditions for the case of intermediate inflation.
2. Power-law inflation with the following Hubble parameter and scale factor
$H(t)=\frac{m}{t},$ (98)
$a(t)=a_{s}t^{m},$ (99)
where $m>1$.
Power-law inflation was considered earlier in large numbers of works (see, for example, in [2, 83], for a review). Also, power-law inflation based on $f(T)$-gravity was considered in [80].
Slow-roll parameters (39)–(40) corresponding to Hubble parameter (98) are
$\epsilon=\delta=\frac{1}{m}=constant.$ (100)
Thus, in this case, there is no exit from inflation by violation of the slow-roll conditions as well.
To obtain the exit from inflation for the case of intermediate and power-law inflation the possibility of implementing the additional reheating process for modified potentials compared to the TEGR case was considered in [80]. However, we can consider the third possible type of inflationary dynamics, which implies the exit from inflation by violation of the slow-roll conditions. This type of exit from inflation corresponds to the similar types of potentials as for the case of TEGR.
3. Inflation with linear deviation from pure de Sitter expansion and following Hubble parameter and scale factor
$H(t)=\lambda-\mu t,$ (101)
$a(t)=a_{s}\exp\left(\lambda t-\frac{\mu}{2}t^{2}\right),$ (102)
where $\lambda>0$ and $\mu>0$ are and some constants.
Slow-roll parameters (39)–(40) corresponding to Hubble parameter (101) are
$\epsilon=\frac{\mu}{H^{2}}=\mu\left(\lambda-\mu t\right)^{-2},~{}~{}~{}~{}\delta=0.$ (103)
Since slow-roll parameter (103) is growing function of the cosmic time, in this case, the end of inflation by violation of the slow-roll conditions $\epsilon_{e}=1$ can be realised for $t_{e}=\frac{\lambda}{\mu}\pm\frac{1}{\sqrt{\mu}}$.
Corresponding condition (90) for $\delta=0$ is
$(1+n)\frac{\tilde{\delta}}{\tilde{\epsilon}}-\frac{n}{2}=0.$ (104)
Thus, we will consider (104) as the condition of the exit from inflation in cosmological models under consideration.
VII Inflationary models with linear dependence $r\sim(1-n_{S})$
Now we will consider the dependence of the tensor-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$. Since the value of the spectral index of scalar perturbations, $n_{S}\simeq 0.97$ and $1-n_{S}\simeq 0.03\ll 1$, we can write the dependence $r=r(1-n_{S})$ as follows,